# Lie Group Theory of Multipole Moments and Shape of Stationary Rotating Fluid Bodies

### v3

@@ 6.1.2 Solving Functional Equations for the Height Function with the Shift Operator
where ${\bm{y}}={\bm{x}}+{\bm{X}}$, and ${\bm{X}}={\bm{X}}({\bm{x}})$ denotes the displacement vector of the level surface. This equation may be directly compared with the condition (1.7) from the Zharkov-Trubitsyn theory [2], which defines the level surfaces of the total potential. The comparison reveals that the Lie group approach identifies $A_{0}(r)\equiv U(r)$, where $U(r)$ is the effective gravitational potential of the reference fluid configuration, assumed to be known. This identification enables a significantly more efficient methodology for determining the shape of equipotential surfaces and the figure of a rotating fluid body. By employing the shift operator technique, this approach offers substantial advantages over the classical Zharkov-Trubitsyn framework [2]. The details of this method are presented in the following sections.
@@
The second method for determining the height function $X$ generalizes the Clairaut equation (3.51) to the nonlinear regime, accounting for finite deformations of the level surfaces. This approach involves a detailed manipulation of the shift operator $\hat{\mathsf{T}}_{\bm{X}}$ to derive a fundamental differential equation for $X({\bm{x}})$ that eliminates the explicit dependence on the gravitational perturbation $K({\bm{x}})$. A comprehensive discussion of both methods is presented in the following two sections.
@@ 6.3.1 Master Equation for the Deformation Gradient
The most advanced developments in this field originate from the Zharkov-Trubitsyn theory of figures [2], which has been extended up to seventh order ($l=7$, corresponding to $\mathsf{m}^{7}$) [9]. However, the feasibility of further extending perturbative methods remains uncertain. Deriving higher-order equations for $X_{l}$ leads to increasingly complex, nonlinear integro-differential equations with a rapidly growing number of terms, necessitating extensive symbolic computation. Moreover, the numerical evaluation of higher-order contributions is prone to instability, as small inaccuracies in lower-order terms can propagate and significantly distort the final results. These challenges underscore the need for carefully designed algorithms and robust numerical strategies [106].
@@
This equation represents a central result in the theory of figures of rotating fluid bodies. Although exact, it is highly nonlinear due to the dependence of the inverse matrix $M^{-1}_{ij}$ on the deformation gradient AijA{ij}, which itself is expressed as an infinite Neumann series – see Eqs. (6.36) and (6.37).
@@ 6.3.3 Master Equation for the Shape Function
The master equation (6.67) for the shape function $f=f({\bm{x}})$ is a new result not previously obtained in the literature. It is exact, depends on all three spatial coordinates, and is nonlinear. Remarkably, in the case of constant density, where the parameter ${\upbeta}=1$, Eq. (6.67) admits several exact solutions. Among these are the classical Maclaurin spheroid and the Jacobi triaxial ellipsoid, both of which are discussed in detail in Appendix 9. The explicit formulation of the master equation for the shape function provides a rigorous foundation for establishing the asymptotic nature of the perturbation series previously employed by various researchers in the study of rotating fluid bodies in both geophysical and astrophysical contexts [107].
@@ 7.1 Spectral Decomposition of the Shape Function
$\rho\frac{dU}{dr}-\frac{dp}{dr}$ $=$ $\frac{4\pi G\rho}{3}\frac{d}{dr}\bigl{[}\bar{\rho}r^{2}(A_{0}-1)\bigr{]}\;,$ (7.5)
@@ 8.5.2 Multipole Moments
$J_{2}$ $=$ $-f_{2}-\frac{\mathsf{m}}{3}-\frac{9\mathsf{m}}{14}f_{2}-\frac{43}{14}f^{2}_{2}\;,$ (8.50)
@@
${\cal M}_{l}$ $=$ $\int_{\mathcal{V}}\left[\rho+\left(-X+XX^{\prime}\right)\rho^{\prime}+\frac{1}{2}X^{2}\rho^{\prime\prime}\right]r^{l+2}P_{l}(\cos\theta)drd^{2}\Omega({\bm{n}})$ (8.52)
$+$ $\oint_{{\mathbb{S}}^{2}}d^{2}\Omega({\bm{n}}_{a})P_{l}\left(\cos\theta_{a}\right)Xa^{l+1}\biggl{[}a\rho+\left(1+\frac{l}{2}\right)\rho X-\frac{1}{2}aX\rho^{\prime}\biggr{]}\;.$ (8.53)
@@ 9.1 Exact Solutions Leading to Maclaurin and Jacobi Ellipsoids
$R^{2}\left[1+\chi(\theta,\varphi)\right]$ $=$ $r\;,$ (9.8)
@@
where $r_{0}$ denotes the radius of the reference configuration, and $R(r_{0},\theta,\varphi)$ describes its deformation. This equation can be recast in matrix form:
@@ 9.2.6 Love Numbers and Gravitational Multipole Moments
To compare with the results of Hubbard [98], we express the external gravitational field in terms of the dimensionless multipole moments $J_{l}$, which are related to ${\cal I}_{l}$ via Eq. (7.61). Hubbard uses a different parameter $\mathfrak{q}$ to characterize the rotational rate, related to $\mathsf{m}$ by Eq. (2.20).
In terms of $\mathfrak{q}$, the multipole moments are expressed as:
@@ Appendix A Relationship between Exponential Flow and Translation Operators
In this appendix, we derive a formula that establishes a connection between the shift of the argument $\bm{x}$} of a smooth function $V({\bm{x}})$ by ${\bm{X}}_{\tau}$ and the exponential map generated by the vector field $\bm{\xi}$}, which induces the shift ${\bm{X}}_{\tau}$. The key identity is:
$\exp\left(\tau L_{\bm{\xi}}\right)V({\bm{x}})$ $=$ $V({\bm{x}}+{\bm{X}}_{\tau})\;,$ (A.1)
where $L_{\bm{\xi}}=\xi^{i}\partial_{i}$ is the Lie derivative along $\bm{\xi}$}, and ${\bm{X}}_{\tau}={\bm{X}}_{\tau}({\bm{x}})$ is defined by Eq. (4.5). This identity can also be expressed in the form:
@@
$\sum_{n=p}^{\infty}\frac{\tau^{n}}{n!}{\bf B}^{i_{1}i_{2}\ldots i_{p}}_{n,p}\left(L_{\bm{\xi}}{\bm{x}},L_{\bm{\xi}}^{2}{\bm{x}},...,L_{\bm{\xi}}^{n-p+1}{\bm{x}}\right)$ $=$ $\frac{1}{p!}X^{i_{1}}_{\tau}X^{i_{2}}_{\tau}\ldots X^{i_{p}}_{\tau}\;.$ (A.10)
@@
$V_{\tau}$ $=$ $\sum_{p=0}^{\infty}\frac{1}{p!}X^{i_{1}}_{\tau}X^{i_{2}}_{\tau}\ldots X^{i_{p}}_{\tau}\,\partial_{i_{1}i_{2}\ldots i_{p}}V({\bm{x}})=V\left({\bm{x}}+{\bm{X}}_{\tau}\right)\;,$ (A.11)
@@ Appendix C Compatibility Conditions for the Master Equation of the Height Function
$n^{q}M^{-1}_{[ip}({\bm{X}})\partial_{p}M^{-1}_{j]q}({\bm{X}})$ $=$ $0\qquad\qquad\Longrightarrow\qquad\qquad\leavevmode\nobreak\ M^{-1}_{[ip}\partial_{p}N_{j]}=0\;,$ (C.2)
@@ Appendix D Calculating the Deformation Gradient Matrix
$N^{i}=M^{-1}_{ij}n^{j}$ $=$ $-\frac{{\cal P}^{ij}\partial_{j}X}{\displaystyle\leavevmode\nobreak\ \Biggl{(}1+X^{\prime}\Biggr{)}\left(1+\frac{X}{r}\right)}+\frac{n^{i}}{1+X^{\prime}}\;.$ (D.17)
@@ E.2 Wigner’s Decomposition of Legendre Polynomial Products
$P_{n_{1}}(\cos\theta)P_{n_{2}}(\cos\theta)P_{n_{3}}(\cos\theta)$ $=$ $\sum_{a_{2}=|n_{2}-a_{1}|}^{n_{2}+a_{1}}\sum_{a_{3}=|n_{3}-a_{2}|}^{n_{3}+a_{2}}T^{0n_{1}a_{1}}T^{a_{1}n_{2}a_{2}}T^{a_{2}n_{3}a_{3}}P_{a_{3}}(\cos\theta)\;,$ (E.16)

### v4

@@ 6.1.2 Solving Functional Equations for the Height Function with the Shift Operator
where ${\bm{y}}={\bm{x}}+{\bm{X}}$, and ${\bm{X}}={\bm{X}}({\bm{x}})$ denotes the displacement vector of the level surface. This equation may be directly compared with the condition (1.7) from the Zharkov-Trubitsyn theory [2], which defines the level surfaces of the total potential. The comparison reveals that the Lie group approach identifies $(4\pi G/3)r^{2}A_{0}(r)\equiv U(r)$, where $U(r)$ is the effective gravitational potential of the reference fluid configuration, assumed to be known. This identification enables a significantly more efficient methodology for determining the shape of equipotential surfaces and the figure of a rotating fluid body. By employing the shift operator technique, this approach offers substantial advantages over the classical Zharkov-Trubitsyn framework [2]. The details of this method are presented in the following sections.
@@
The second method for determining the height function $X$ generalizes the Molodensky equation (3.49) to the nonlinear regime, accounting for finite deformations of the level surfaces. This approach involves a detailed manipulation of the shift operator $\hat{\mathsf{T}}_{\bm{X}}$ to derive a fundamental differential equation for $X({\bm{x}})$ that eliminates the explicit dependence on the gravitational perturbation $K({\bm{x}})$. A comprehensive discussion of both methods is presented in the following two sections.
@@ 6.3.1 Master Equation for the Deformation Gradient
The most advanced developments in this field originate from the Zharkov-Trubitsyn theory of figures [2], which has been extended up to seventh order ($l=14$, corresponding to $\mathsf{m}^{7}$) [9]. However, the feasibility of further extending perturbative methods remains uncertain. Deriving higher-order equations for $X_{l}$ leads to increasingly complex, nonlinear integro-differential equations with a rapidly growing number of terms, necessitating extensive symbolic computation. Moreover, the numerical evaluation of higher-order contributions is prone to instability, as small inaccuracies in lower-order terms can propagate and significantly distort the final results. These challenges underscore the need for carefully designed algorithms and robust numerical strategies [106].
@@
This equation represents a central result in the theory of figures of rotating fluid bodies. Although exact, it is highly nonlinear due to the dependence of the inverse matrix $M^{-1}_{ij}$ on the deformation gradient $A_{ij}$, which itself is expressed as an infinite Neumann series – see Eqs. (6.36) and (6.37).
@@ 6.3.3 Master Equation for the Shape Function
The master equation (6.67) for the shape function $f=f({\bm{x}})$ is a new result not previously obtained in the literature. It is exact, depends on all three spatial coordinates, and is nonlinear. Remarkably, in the case of constant density, where the parameter ${\upbeta}=1$, Eq. (6.67) admits several exact solutions. Among these are the classical Maclaurin spheroid and the Jacobi triaxial ellipsoid, both of which are discussed in detail in Section 9. The explicit formulation of the master equation for the shape function provides a rigorous foundation for establishing the asymptotic nature of the perturbation series previously employed by various researchers in the study of rotating fluid bodies in both geophysical and astrophysical contexts [107].
@@ 7.1 Spectral Decomposition of the Shape Function
$\rho\frac{dU}{dr}-\frac{dp}{dr}$ $=$ $-\frac{4\pi G\rho}{3}\frac{d}{dr}\left(r^{2}A_{0}-\frac{3U}{4\pi G}\right)\;,$ (7.5)
@@ 8.5.2 Multipole Moments
$J_{2}$ $=$ $-f_{2}-\frac{\mathsf{m}}{3}-\frac{\mathsf{m}}{7}f_{2}-\frac{11}{7}f^{2}_{2}\;,$ (8.50)
@@
${\cal M}_{l}$ $=$ $\int_{\mathcal{V}}\left[\rho+\left(-X+XX^{\prime}\right)\rho^{\prime}+\frac{1}{2}X^{2}\rho^{\prime\prime}\right]r^{l+2}P_{l}(\cos\theta)drd^{2}\Omega({\bm{n}})$
$+$ $\oint_{{\mathbb{S}}^{2}}d^{2}\Omega({\bm{n}}_{a})P_{l}\left(\cos\theta_{a}\right)Xa^{l+1}\biggl[a\rho+\left(1+\frac{l}{2}\right)\rho X-\frac{1}{2}aX\rho^{\prime}\biggr]\;.$
@@ 9.1 Exact Solutions Leading to Maclaurin and Jacobi Ellipsoids
$R^{2}\left[1+\chi(\theta,\varphi)\right]$ $=$ $r^{2}\;,$ (9.8)
@@
where $r_{0}$ denotes the radius of the reference configuration, and $R\equiv R(r_{0},\theta,\varphi)$ describes its deformation. This equation can be recast in matrix form:
@@ 9.2.6 Love Numbers and Gravitational Multipole Moments
To compare with the results of Hubbard [98], we express the external gravitational field in terms of the dimensionless multipole moments $J_{l}$, which are related to ${\cal I}_{l}$ via Eq. (7.61). Hubbard uses a different parameter $\mathsf{q}$ to characterize the rotational rate, related to $\mathsf{m}$ by Eq. (2.20).
In terms of $\mathsf{q}$, the multipole moments are expressed as:
@@ Appendix A Relationship between Exponential Flow and Translation Operators
In this appendix, we derive a formula that establishes a connection between the shift of the argument $\bm{x}$} of a smooth function $V({\bm{x}})$ by ${\boldsymbol{\mathscr{X}}}_{\tau}$ and the exponential map generated by the vector field $\bm{\xi}$}, which induces the shift ${\boldsymbol{\mathscr{X}}}_{\tau}$. The key identity is:
$\exp\left(\tau L_{\bm{\xi}}\right)V({\bm{x}})$ $=$ $V({\bm{x}}+{\boldsymbol{\mathscr{X}}}_{\tau})\;,$ (A.1)
where $L_{\bm{\xi}}=\xi^{i}\partial_{i}$ is the Lie derivative along $\bm{\xi}$}, and ${\boldsymbol{\mathscr{X}}}_{\tau}={\boldsymbol{\mathscr{X}}}_{\tau}({\bm{x}})$ is defined by Eq. (4.5). This identity can also be expressed in the form:
@@
$\sum_{n=p}^{\infty}\frac{\tau^{n}}{n!}{\bf B}^{i_{1}i_{2}\ldots i_{p}}_{n,p}\left(L_{\bm{\xi}}{\bm{x}},L_{\bm{\xi}}^{2}{\bm{x}},...,L_{\bm{\xi}}^{n-p+1}{\bm{x}}\right)$ $=$ $\frac{1}{p!}{\mathscr{X}}^{i_{1}}_{\tau}{\mathscr{X}}^{i_{2}}_{\tau}\ldots{\mathscr{X}}^{i_{p}}_{\tau}\;.$ (A.10)
@@
$V_{\tau}$ $=$ $\sum_{p=0}^{\infty}\frac{1}{p!}{\mathscr{X}}^{i_{1}}_{\tau}{\mathscr{X}}^{i_{2}}_{\tau}\ldots{\mathscr{X}}^{i_{p}}_{\tau}\,\partial_{i_{1}i_{2}\ldots i_{p}}V({\bm{x}})=V\left({\bm{x}}+{\boldsymbol{\mathscr{X}}}_{\tau}\right)\;,$ (A.11)
@@ Appendix C Compatibility Conditions for the Master Equation of the Height Function
$n^{q}M^{-1}_{[ip}({\bm{X}})\partial_{p}M^{-1}_{j]q}({\bm{X}})$ $=$ $0\qquad\qquad\Longrightarrow\qquad\qquad\penalty 10000\ M^{-1}_{[ip}\partial_{p}N_{j]}=0\;,$ (C.2)
@@ Appendix D Calculating the Deformation Gradient Matrix
$N^{i}=M^{-1}_{ij}n^{j}$ $=$ $-\frac{{\cal P}^{ij}\partial_{j}X}{\displaystyle\penalty 10000\ \Biggl(1+X^{\prime}\Biggr)\left(1+\frac{X}{r}\right)}+\frac{n^{i}}{1+X^{\prime}}\;.$ (D.17)
@@ E.2 Wigner’s Decomposition of Legendre Polynomial Products
$P_{n_{1}}(\cos\theta)P_{n_{2}}(\cos\theta)P_{n_{3}}(\cos\theta)$ $=$ $\sum_{a_{2}=|n_{2}-n_{1}|}^{n_{2}+n_{1}}\sum_{a_{3}=|n_{3}-a_{2}|}^{n_{3}+a_{2}}T^{n_{1}n_{2}a_{2}}T^{a_{2}n_{3}a_{3}}P_{a_{3}}(\cos\theta)\;,$ (E.16)